OpenStudy (princessaurora):

** POR FAVOR WILL FAN AND MEDAL** Situation: You invest $900 in an account that pays an interest rate of 5.5%, compounded continuously. Calculate the balance of your account after 6 years.

OpenStudy (princessaurora):

They gave me a formula but it's not working I think there's a glitch and it won't show me the rest of it I think

OpenStudy (mathmate):

The formula for continuous interest problems is \(FV=PVe^{rt}\) FV=future value PV=present value r=annual rate of interest (in decimal, not percent) t=time in years e=constant, 2.71828....

OpenStudy (princessaurora):

Is the future value 6 years?

OpenStudy (princessaurora):

Please help me set it up I'm getting confused.

OpenStudy (princessaurora):

6=(900)(2.71828)^(5.5)(6) ?

OpenStudy (fortytherapper):

You have FV at 6, but you're looking for FV

OpenStudy (mathmate):

yes, time t=6 (years) PV is the current investment FV is what you expect to get after t=6 years r=rate of interest=5.5%=0.055 in decimal

OpenStudy (princessaurora):

So, FV=(900)(2.71828)^(0.055)(6) ?

OpenStudy (mathmate):

Note that interest is in decimal form 0.055 and there are parentheses between r and t, i.e. the exponent is for the product r*t. The way you interpreted the formula will have the time t multiplied (instead of exponentiated).

OpenStudy (princessaurora):

can you write it out please I keep writing it wrong?

OpenStudy (mathmate):

FV=PV*e^(rt) is what you need to put on your calculator.

OpenStudy (princessaurora):

I know but you said what I set up is wrong

OpenStudy (mathmate):

Read up PEMDAS to help you understand how formulas work: http://www.purplemath.com/modules/orderops.htm

OpenStudy (mathmate):

You may have noticed that there are parentheses around r and t. Without the parentheses, only r will be exponentiated.

OpenStudy (princessaurora):

oh okay I see!

OpenStudy (princessaurora):

I got 1251.87

OpenStudy (princessaurora):

is that worng

OpenStudy (fortytherapper):

Nope, I got the same thing. Good one

OpenStudy (princessaurora):

I GOT IT RIGHT Y'ALL THANK THE HEAVENS

OpenStudy (fortytherapper):

Yay!

OpenStudy (princessaurora):

@FortyTheRapper Thanks for your help :)

OpenStudy (fortytherapper):

Thanks but I give all credit to @mathmale on this one